Suppose that the individual demand functions for a particular good are Q 30- 4P for 4 people and Q 20- P for 3 people. Assume that these 7 people make up the entire market and act as price-takers. The marginal cost of producing the good is constant at MC-$5.

(a) If the good described above is delicious ice cream sundaes, what is the Pareto efficient price and quantity in this market? (b) If the good is instead fireworks, then what is the Pareto efficient quantity in this market? Hint: Draw the graph and pay particular attention to the SMB curve
Answer #1

a) Given price taker. Thus perfectly competitive market. P=MC . Qi=30-4*5=10. Since 4 people total quantity is 4*10=40

Qi=20-5=15. Since 3 people , we have quantity 45. Thus total quantity is 85

b) Since fireworks can be enjoyed by all individuals even if one person buys, Thus price is 5. quantity is 10